# YEAST: Yet Another Sequential Test

### v2

@@ Abstract
Online evaluation of machine learning models is typically conducted through A/B experiments. Sequential statistical tests are valuable tools for analysing these experiments, as they enable researchers to stop data collection early without increasing the risk of false discoveries. However, existing sequential tests either limit the number of interim analyses or suffer from low statistical power. In this paper, we introduce a novel sequential test designed for continuous monitoring of online experiments. We validate our method using semi-synthetic simulations and demonstrate that it outperforms current state-of-the-art sequential testing approaches. Our method is derived using a new technique that “inverts” a bound on the probability of threshold crossing, based on a classical maximal inequality.
@@ 1 Introduction
Online A/B experiments have become a standard approach for evaluating the impact of machine learning models on the systems and products where they are deployed. While this approach offers significant advantages, it also poses the risk of exposing users to harmful changes that could degrade their experience. Unlike traditional techniques like the Student t-test, sequential testing supports repeated significance checks without increasing the false positive rate. This makes it well-suited for real-time monitoring of experiments, allowing early detection of harmful effects, which helps protect user experience and reduce financial losses.
@@ 3.3.2 Setting the Input Parameters
While the need to set a finite horizon $N$ can be viewed as a disadvantage of YEAST, we have not found it problematic in practice. When running A/B experiments it is typical to plan for a certain duration and the number of observations collected during the experiment time frame is usually predictable. In addition, the decision boundary of YEAST depends on the square root of N, which makes the method more robust to inaccuracies in the estimation of N. That said, we fully acknowledge that there are situations where it is difficult to reliably estimate $N$ and in that case, it may indeed be better to resort to other methods (e.g. [11]) that do not require setting a finite horizon.
In the empirical evaluations in Section 4, we never assume that $N$ is known and estimate it from pre-experiment data (as part of the method). Hence, the respective empirical results incorporate the uncertainty in the estimation of $N$. The estimation was done by simply setting $N$ to the number of events collected during a pre-experiment period of the same length as the duration of the experiment.
@@ 4 Validation Using Real-World Data: a Semi-Synthetic Experiment
LanDeMetsOBF The GST method [8] with the O’Brien-Fleming alpha-spending function [23]. For computation reasons, we constructed the boundary using 100 interim checkpoints and then extended it in a piecewise-constant manner to support continuous monitoring.
@@ 6 Conclusion
The authors are grateful to Evan Miller for insightful comments and his excellent blog [22], which sparked the idea for this research.
@@ B.1 Simulation Results
LanDeMetsOBF The GST method [8] with the O’Brien-Fleming alpha-spending function [23]. The method was applied in the continuous mode (i.e. the interim analyses were performed after each observation).
@@
All of the compared methods were employed in the continuous monitoring mode meaning that the check for significance was performed after each observation. In Appendix C we report additional evaluation results for the case where the methods are used in the "discrete mode" (i. e. with a fixed number of interim significance checks).
For each experimental setting, we conducted 100,000 replications. Each replication can result in a detection or no-detection. A detection occurs when the respective test flags significance (at any point of the monitoring process). We compute the share of replications where a detection occurs. When the treatment does not have an effect this share is the so-called (empirical) false detection rate (or, synonymously, false positive rate, type-I error, or test “size”). In settings where the treatment has an effect, this share is the (empirical) power. We report the measured false detection rate and power in Table 4.
@@
1 YEAST 0.05 0.45 0.92 1.00 1.00
2 YEASTnv110 0.04 0.40 0.90 1.00 1.00
3 YEASTnv120 0.03 0.36 0.88 1.00 1.00
4 YEASTnv80 0.07 0.52 0.93 1.00 1.00
5 YEASTnv90 0.06 0.45 0.93 1.00 1.00
6 mSPRTphi100 0.02 0.22 0.76 0.99 1.00
7 mSPRTphi11 0.03 0.22 0.73 0.98 1.00
8 mSPRTphi25 0.03 0.24 0.76 0.99 1.00
9 GAVI250 0.02 0.24 0.76 0.99 1.00
10 GAVI500 0.02 0.23 0.76 0.99 1.00
11 GAVI750 0.01 0.21 0.75 0.99 1.00
12 Bonferroni 0.00 0.04 0.40 0.90 0.99
13 LanDeMetsOBF 0.06 0.46 0.93 1.00 1.00
14 SeqC2ST-QDA 0.02 0.04 0.09 0.22 0.41

### v3

@@ Abstract
Online evaluation of machine learning models is typically conducted through A/B experiments. Sequential statistical tests are valuable tools for analysing these experiments, as they enable researchers to stop data collection early without increasing the risk of false discoveries. However, existing sequential tests either limit the number of interim analyses or suffer from low statistical power. In this paper, we introduce a novel sequential test designed for continuous monitoring of A/B experiments. We validate our method using semi-synthetic simulations and demonstrate that it outperforms current state-of-the-art sequential testing approaches. Our method is derived using a new technique that “inverts” a bound on the probability of threshold crossing, based on a classical maximal inequality.
@@ 1 Introduction
Online A/B experiments have become a standard approach for evaluating the impact of machine learning models on the systems and products where they are deployed. While this approach offers significant advantages, it also poses the risk of exposing users to harmful changes that could degrade their experience. Unlike traditional techniques such as Student’s t test, sequential testing supports repeated significance checks without increasing the false positive rate. This makes it well-suited for real-time monitoring of experiments, allowing early detection of harmful effects, which helps protect user experience and reduce financial losses.
@@ 3.3.2 Setting the Input Parameters
While the need to set a finite horizon $N$ can be viewed as a disadvantage of YEAST, we have not found it problematic in practice. When running A/B experiments it is typical to plan for a certain duration and the number of observations collected during the experiment time frame is usually predictable. In addition, the decision boundary of YEAST depends on the square root of $N$, which makes the method more robust to inaccuracies in the estimation of $N$. That said, we acknowledge that there are situations where it is difficult to reliably estimate $N$ and in that case, it may indeed be better to resort to other methods (e.g. [11]) that do not require setting a finite horizon. In the empirical evaluations in Section 4, we never assume that $N$ is known and estimate it from pre-experiment data (as part of the method). Hence, the respective empirical results incorporate the uncertainty in the estimation of $N$.
@@ 4 Validation Using Real-World Data: a Semi-Synthetic Experiment
LanDeMetsOBF The GST method [8] with the O’Brien-Fleming alpha-spending function [23]. For computational reasons, we constructed the boundary using 100 interim checkpoints and then extended it in a piecewise-constant manner to support continuous monitoring.
@@ 6 Conclusion
The authors are very grateful to Evan Miller for insightful comments and his excellent blog [22], which sparked the idea for this research.
@@ B.1 Simulation Results
LanDeMetsOBF The GST method [8] with the O’Brien-Fleming alpha-spending function [23]. For computational reasons, we constructed the boundary using 100 interim checkpoints and then extended it in a piecewise-constant manner to support continuous monitoring.
@@
All of the compared methods were employed in the continuous monitoring mode meaning that the check for significance was performed after each observation. In Appendix C we report additional evaluation results for the case where YEAST was employed in the "discrete mode" (i. e. with a fixed number of interim significance checks).
For each experimental setting, we conducted 100,000 replications. Each replication can result in a detection or no-detection. A detection occurs when the respective test flags significance (at any point of the monitoring process). We compute the share of replications where a detection occurs. When the treatment does not have an effect this share is the so-called (empirical) false detection rate (or, synonymously, false positive rate, type-I error, or test “size”). In settings where the treatment has an effect, this share is the (empirical) power. Table 4 presents the measured false detection rate and power, along with their corresponding 95% confidence intervals adjusted for multiple comparisons.
@@
1 YEAST 0.047 ± 0.002 0.448 ± 0.005 0.902 ± 0.003 0.995 ± 0.001
2 YEASTn110 0.038 ± 0.002 0.408 ± 0.005 0.883 ± 0.003 0.994 ± 0.001
3 YEASTn120 0.030 ± 0.002 0.372 ± 0.004 0.864 ± 0.003 0.992 ± 0.001
4 YEASTn80 0.075 ± 0.002 0.534 ± 0.005 0.933 ± 0.002 0.997 ± 0.000
5 YEASTn90 0.059 ± 0.002 0.490 ± 0.005 0.918 ± 0.003 0.996 ± 0.001
6 mSPRT100 0.016 ± 0.001 0.235 ± 0.004 0.751 ± 0.004 0.976 ± 0.001
7 mSPRT011 0.032 ± 0.002 0.249 ± 0.004 0.739 ± 0.004 0.972 ± 0.002
8 mSPRT025 0.028 ± 0.002 0.260 ± 0.004 0.758 ± 0.004 0.976 ± 0.001
9 GAVI250 0.025 ± 0.001 0.260 ± 0.004 0.763 ± 0.004 0.977 ± 0.001
10 GAVI500 0.019 ± 0.001 0.245 ± 0.004 0.758 ± 0.004 0.977 ± 0.001
11 GAVI750 0.014 ± 0.001 0.227 ± 0.004 0.744 ± 0.004 0.975 ± 0.001
12 Bonferroni 0.008 ± 0.001 0.067 ± 0.002 0.440 ± 0.005 0.877 ± 0.003
13 LanDeMetsOBF 0.054 ± 0.002 0.466 ± 0.005 0.927 ± 0.002 0.999 ± 0.000
14 SeqC2ST_QDA 0.022 ± 0.001 0.037 ± 0.002 0.092 ± 0.003 0.220 ± 0.004